OpenStudy (amonoconnor):

Is this the right answer? "Approximate: sqrt.(99.8)" Any and all help is greatly appreciated!

OpenStudy (freckles):

you know sqrt(100)

OpenStudy (amonoconnor):

Yes! :)

OpenStudy (freckles):

so I suggest doing linerzation for f(x)=sqrt(x) at x=100

OpenStudy (anonymous):

9.98

OpenStudy (freckles):

first step to do this is find the tangent line at x=100 \[y-f(100)=f'(100)(x-100) \\ y=f'(100)(x-100)+f(100)\]

OpenStudy (freckles):

this equation could be a pretty close approximation for values near x=100 and 99.8 is pretty close to 100

OpenStudy (amonoconnor):

Right, that's what I did, I just wanted to make sure I did it right. It's been a bit lols;) I got 9.99 with the Linearization formula dealio. Is that close enough to the actual answer?

OpenStudy (freckles):

\[\sqrt{x} \approx f'(100)(x-100)+f(100) \text{ for values near } x=100\]

OpenStudy (freckles):

yep it should be in that ball park it should be something pretty close to 10 which is what you got

OpenStudy (freckles):

but I'm going to see what I get using the equation I have above

OpenStudy (freckles):

yep I got 999/100 as well which is 9.99

OpenStudy (amonoconnor):

Sweet!! Thanks a zillion;)

OpenStudy (freckles):

putting it into the calculator gives me 9.989994... which is pretty damn close to what we got

OpenStudy (amonoconnor):

Heck yes it is... sounds good to me:)